OpenStudy (anonymous):

96=-6(-4-3y) 96=24-3y 96=24-3y +24 +24 120=3y 120/3=40 3y/3=3 Obviously I messed something up because 40 is not one of the answer choices.. A) y=2 B) y=4 C) y=-2 D) y=-4

OpenStudy (ranga):

a(b+c) = ab + ac -6(-4-3y) = 24 + 18y

OpenStudy (anonymous):

24+18y=96 -24 -24 18y 72 18y/18=18 72/18=4 Now I got it! :D

OpenStudy (ranga):

Yep! Congrats!

OpenStudy (anonymous):

Thank you, it was all thanks to you.

OpenStudy (ranga):

u r welcome.

OpenStudy (mathstudent55):

\(96=\color{red}{-6}(-4\color{blue}{-3y})\) \(96=24 \color{red}{-6} (\color{blue}{- 3y})\) \(96=24 + 18y \)